The prices of tickets for the local rock concert were $16 for Friday night and $28 for Saturday night. The total attendance for the two nights was 1090. If the total receipts were $24,232, how many people attended Saturday night? Let X = number of people that attended Friday’s concert and Y = number of people that attended Saturday’s concert.
I dont understand how to solve it.

Answer by checkley71(8403)

X+Y=1090 & 16X+28Y=24232 OR X=1090-Y SUBSTITUTING (1090-Y) FOR X IN EQUATION #2 WE GET
16(1090-Y)=24232 OR 17440-16Y+28Y=24232 OR 12Y=6792 0R Y=6792/12 OR Y=566 THUS
1090-566=524 OR FRI. SALES WERE 524@16=8384 & SAT SALES WERE 566@28=15848 SO
8384+15848=24232